A nutritionist wants to know the population proportion of grade 11 students who eat vegetables . Pegged at a confidence of 95% survey among 1200 respondents was conducted and 200 said that they eat vegetables. What is the margin of error???

1
Expert's answer
2022-06-16T10:03:06-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

"ME=z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p}}{n}}""\hat{p}=\dfrac{x}{n}=\dfrac{200}{1200}=\dfrac{1}{6}""ME=1.96\times\sqrt{\dfrac{\dfrac{1}{6}(1-\dfrac{1}{6})}{1200}}\approx0.021"